OpenStudy (anonymous):

Find the exact circumference of a circle with the given radius. 3¼ inches C = 6.25π in. 6.5π in. 23.25π in.

Vocaloid (vocaloid):

circumference = 2*pi*r

OpenStudy (theraggedydoctor):

What is the formula for finding circumference of a circle?

OpenStudy (anonymous):

c=pi x r^2

OpenStudy (anonymous):

I think

Vocaloid (vocaloid):

nope, pi(r^2) is area, not circumference circumference = 2*pi*r

OpenStudy (anonymous):

oh haha thanks. How would you solve that?

Vocaloid (vocaloid):

radius = 3 1/4 inches circumference = 2*pi*radius plug in radius and multiply

OpenStudy (anonymous):

thanks!